OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of .

OpenStudy (lgbasallote):

slope of what?

OpenStudy (anonymous):

with a slope of -3

OpenStudy (lgbasallote):

use the formula \[y - y_1 = m(x-x_1)\] where m = slope y1 = y coordinate x1 = x coordinate does that help?

OpenStudy (anonymous):

y1+4 = -3(x1-1)

OpenStudy (anonymous):

the choices are y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1 and i think its c? am i right?

OpenStudy (lgbasallote):

hmm did you write the question right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

hmm weird...none of the choices you gave is the answer to the question you gave

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (lgbasallote):

lol..you gave the wrong question

OpenStudy (lgbasallote):

C is not correct sorry

OpenStudy (anonymous):

It would y=-3x-1

OpenStudy (lgbasallote):

using the formula i gave this will give you \[y + 5 = -3(x-2)\] distribute \[\implies y + 5= - 3x + 6\] subtract 5 from both sides \[\implies y = -3x + 6 - 5\] combine \[\implies y = -3x + 1\] got it @danireynoso ?

OpenStudy (anonymous):

thanks a ton!